OpenStudy (anonymous):

Factor the algebraic expression x2 - z4

OpenStudy (anonymous):

\[x^2 - z^4\]

OpenStudy (cggurumanjunath):

(x)^2-(z^2)^2

OpenStudy (cggurumanjunath):

(x+z^2)(x-z^2)

OpenStudy (anonymous):

Jeez you're good :D Thank you so much

OpenStudy (cggurumanjunath):

(a+b)(a-b)=(a^2-b^2) a=x b=z^2

OpenStudy (cggurumanjunath):

Anytime.

OpenStudy (cggurumanjunath):

:)